3. Use the below tables to complete lab calculations on the worksheet on pages 2-6 of this document. You will submit the workWorksheet for Enzyme Kinetics Lab I 1. Complete the table (1 point) Note: All volumes are in mL Buffer Substrate 0.5 0.25 0.1180-120 Average A Abs (475 nm) /min 4. Determine the A Abs (475 nm)/min for the 0.25 mM, 0.5 mm, 1 mM and 0.2 mm substrate re6. Convert each of the molar concentrations determined in number 5 to standard enzyme concentration based rates (umoles dopac9. Determine the Km and Vmax values of Tyrosinase from the graph. (4 points) X-intercept: Y-intercept: Km= V maxWorksheet for Enzyme Kinetics Lab | Name: SAMPLE CALCULATION GUIDE Date: 1. Table 1. Complete the table from page 41. (1 poin4. Using the Lambert-Beer equation, calculate the molar concentration of dopachrome represented by each A Abs (475 nm/min for7. Graph a Lineweaver-Burke plot (1/V. (umal product/min) versus 1/(S) mm) with proper techniques (6 pts) Title: dineweover B
Answer #1

1.

Buffer Volume (mL)

0.96

1.21

1.36

1.41

Substrate Volume (mL)

0.5

0.25

0.1

0.05

Enzyme Volume (mL)

0.04

0.04

0.04

0.04

Total Reaction Volume (mL)

1.5

1.5

1.5

1.5

2.

Time (sec)

Abs 475nm

Substrate 5mM

Abs 475nm

Substrate 2.5mM

Abs 475nm

Substrate 1mM

Abs 475nm

Substrate 0.5mM

30

0.14

0.085

0.078

0.063

60

0.259

0.204

0.145

0.112

90

0.371

0.309

0.205

0.158

120

0.473

0.407

0.262

0.198

150

0.57

0.498

0.313

0.236

180

0.654

0.578

0.359

0.27

3.

Time (sec)

Abs 475nm

Substrate 5mM

Abs 475nm

Substrate 2.5mM

Abs 475nm

Substrate 1mM

Abs 475nm

Substrate 0.5mM

90-30

0.231

0.224

0.127

0.095

120-60

0.214

0.203

0.117

0.086

150-90

0.199

0.189

0.108

0.078

180-120

0.181

0.171

0.097

0.072

Average ∆Abs(475nm)/min

0.206

0.197

0.112

0.083

4.

a. ∆Abs(475nm)/min (5 mM Substrate) = 0.206

b. ∆Abs(475nm)/min (2.5 mM Substrate) = 0.197

c. ∆Abs(475nm)/min (1 mM Substrate) = 0.112

d. ∆Abs(475nm)/min (0.5 mM Substrate) = 0.083

5. A = e C l   C = A /e = A/3600

a. [dopachrome] for 5 mM substrate reaction = 5.73 x 10-5 M/min

b. [dopachrome] for 2.5 mM substrate reaction = 5.47 x 10-5 M/min

c. [dopachrome] for 1 mM substrate reaction = 3.12 x 10-5 M/min

d. [dopachrome] for 0.5 mM substrate reaction = 2.3 x 10-5 M/min

6. mole = molar concentration x total volume of enzyme reaction mixture

a. Using 5 mM substrate:

Moles dopachrome produced per min = 3.82 x 10-8 moles/min

U moles dopachrome produced per min = 0.038 µmoles/min

b. Using 2.5 mM substrate:

Moles dopachrome produced per min = 3.65 x 10-8 moles/min

U moles dopachrome produced per min = 0.036 µmole/min

c. Using 1 mM substrate:

Moles dopachrome produced per min = 2.08 x 10-8 moles/min

U moles dopachrome produced per min =0.021 µmole/min

d. Using 0.5 mM substrate:

Moles dopachrome produced per min = 1.53 x 10-8 moles/min

U moles dopachrome produced per min = 0.015 µmole/min

7.

Vo

[S]

1/ Vo

1/[S]

µmole/min

mM

0.038

5.0

26.31579

0.2

0.036

2.5

27.77778

0.4

0.021

1.0

47.61905

1

0.015

0.5

66.66667

2

8.

y = 23.361x + 21.071 R2 = 0.9828 1N. (umol/min) ........ Linear (0) - 0.5 2 2.5 1 1.5 1/[S] MM

9. Lineweaver-Burk plot    1/V = Km/Vmax [1/S] + 1/Vmax

y-intercept = 1/Vmax = 21.07       Vmax = 0.0475 µmol/min

X-intercept = slope = 23.361

Km/Vmax    = 23.361

Km= Vmax x 23.361 =    0.0475 x 23.361 = 1.1096 mM     

Note : please give your feedback